OpenStudy (anonymous):

|p1f1-pf2|=k=|p1f1-p1f2| find the equation of the hyperbola with foci (+/-7, 0) such that the distance from any point on the hyperbola to the foci differ by 10

OpenStudy (mertsj):

So 2a = 10 and a = 5 c is given to be 7 a^2+b^5=c^2 so you can find b. If you know a and b you can write the equation.

OpenStudy (anonymous):

\[plot (3,-3\pi/4) find the other three additional points using -2\pi <\theta <2\pi \]